# Juliana Xavier

### v1

@@ 1. Introduction
(1) $\limsup\frac{1}{n}\log(\#\operatorname{Fix}(f^{n}))\geq\log d\\$
for a dynamical system $f$ of degree $d$ is both simple and natural. For instance, if $f:S^{1}\to S^{1}$ is continuous with degree $d$, the inequality is easily seen to be true. Indeed, the graph of a lift $\tilde{f}$$\mathbb{R}$$\to{\mathbb{R}}$ must intersect at least d1d-1 times the set $\{y=x+n,n\in{\mathbb{Z}}\}$. This gives at least $d^{n}-1$ fixed points for the map $f^{n}$.
However, in contrast, one can easily construct a continuous map $f:S^{2}\to S^{2}$ with degree $d$ where inequality (1) does not hold. The map expressed in polar coordinates in ${\mathbb{R}}^{2}$ as $(r,\theta)\to(dr,d\theta)$ and extended to the sphere with $\infty\to\infty$, has degree $d$ and just two periodic points.
More interestingly, to this day there is no list of well understood obstructions for the inequality to hold for the two dimensional sphere. Moreover, in trying to undersand this simple yet fundamental question of low dimensional dynamics, there are significant other open questions and problems that arise.
This paper is an introductory survey of basic results to date on this problem, and we also present a number of such open questions and problems.
To finish this introductory section I would like to add that it was Mike Shub that grew awareness around the growth rate inequality problem within the endomorphism group working in Montevideo, Uruguay, during one of his many stays in our country. We are deeply thankful for the many years of both joy and challenge in our work.
@@ 2. $C^{1}$ scenario
We have already seen that if $f$ is a continuous map, the inequality (1) may not hold. On the other hand, if $f$ is a rational map, then the inequality holds (see Proposition 1 of [32]). From [33] it follows that if $f$ is $C^{1}$, then it must have infinitely many distinct periodic points, but their growth rate remains unknown. Of course this rises the question as to whether certain regularity will imply that inequality (1) holds. This is not the focus of the present paper. Nonetheless, we present in this section a very brief overview of the results regarding differentiable maps.
@@
The latitude preserving hypothesis mean that $f$ carries each latitude into another latitude or to one of the poles. It need not be a homeomorphism from one latitude to another. Of course, inequality (1) follows from Theorem 2.1.
@@
These sets of hypothesis (smooth and preserves some “geographical” (singular) foliation ) were later considered in [11], where they considered what happens when one changes the “geography” and replace the assumption of smoothness by another one. More precisely, they specify certain foliations with only one sigularity and ask the map to be smooth near this “pole” therefore obtaining the growth rate inequality.
Their research carried on by studying “toy models” which may help one to understand the mechanism of creating periodic points under the assumption of smoothness. See [12], where they confirm Shub’s conjecture for smooth maps of $S^{2}$ preserving the longitudinal foliation.
These strong hypothesis of preserving a foliation and being of class $C^{1}$ were dropped and replaced by an assumption on the homological action of the map that generalizes the former series of results in [29]:
@@
Let $f:S^{2}\to S^{2}$ be a continuous map such that $\deg f=d,|d|>1$. Assume the following hypothesis hold:
@@
Let $f:S^{2}\to S^{2}$ be a degree $d$ $C^{1}$ endomorphism with $|d|>1$. If $f^{-1}(\{N,S\})\subset\{N,S\}$, then $f$ has at least |d1||d-1| fixed points.
A crucial remark, that is responsible for these last results and that will be exploited in the more detailed discussion to follow, is that the $C^{1}$ hypothesis is only used to guarantee that the map is a sink around its critical points. It is natural then not to impose differentiability, but rather try and understand what the real obstructions are for a sphere endomorphism to verify inequality (1).
@@
[16] Let $f:S^{2}\to S^{2}$ be a degree $d$ map such that all its periodic orbits are isolated as invariant sets. Then, if f has no sources of degree $r$ with $|r|>1$ (this happens, in particular, if $f$ is $C^{1}$) we have that equation (1) holds for $f$.
When a sphere endomorphism $f$ does verify the inequality (1), we say that $f$ has the rate.
@@
Sphere endomorphisms are too large of a universe. It is natural to try and focus on the simplest cases at the beginning. Other than homeomorphisms, the simplest maps one can think of, in terms of their local properties, are covering maps. A covering map $f:X\to Y$ is a continuous map such that every $y\in Y$ has an open neighborhood $V$ whose preimage is a disjoint union of open sets each of one is mapped homeomorphically onto $V$. If one is interested in dynamics though, one needs $X$ to be equal to $Y$, that is, one wants a self-covering $f:X\to X$ in order to be able to iterate. However, the only possible covering maps $f:S^{2}\to S^{2}$ are homeomorphisms. So, if one wants to escape the invertible world, one needs to allow critical points to be present. The best generalization is the following:
@@
We will show in this section that any continuous map of the closed annulus of degree $d$ ($|d|>1$) is semi-conjugate to $z^{d}$ on $S^{1}$. Moreover, if $f:A\to A$ has degree $d$ and $K\subset A$ is an invariant compact essential continuum, then $f:|_{K}:K\to K$ is semi-conjugate to $z^{d}$ on $S^{1}$. However, this is not true for maps acting on the open annulus $A$ (as the ones arising from restricting a branched covering of the sphere with a set of two points that is totally invariant); even if they are covering maps. We refer the reader to Section 7.3 for an example of $f:A\to A$ a covering map of degree $d$ that is not semi-conjugate to $z^{d}$ acting on $S^{1}$. We denote $\tilde{A}={\mathbb{R}}\times(0,1)$ the universal cover of $A$.
@@
Another simple idea gives the rate for maps of the annulus of degree $d$, $|d|>1$: any lift $F$ will push vertical lines that are sufficiently far to the right or left striclty to the right of themselves, or striclty to the left of themselves (we are using $|d|>1$ here). So, with hypothesis guaranteeing that horizontal lines are also pushed strictly above them, or strictly below them, one can calculate the Lefschetz indez of a simple closed curve formed by intersecting appropriate pairs of vertical and horizontal lines. This is formalized as follows:
@@
[19](“Sarkovskii-Iglesias”). Let $F$ be a map of the open annulus $A$ that interchanges the ends of $A$. If $|\deg(F)|>1$, then $F$ has the rate.
@@
The orbifold associated to a Thurston map $f$ is the pair $(\mathcal{O}_{f},\nu_{f})$. Note that $\{p\in S^{2}:\nu_{f}(p)\geq 2\}=P_{f}$ is a finite set. If we label these points $p_{1},\dots,p_{n}$ such that $2\leq\nu_{f}(p_{1})\leq\ldots\leq\nu_{f}(p_{n})$, then the $n$-tuple $(\nu_{f}(p_{1}),\ldots,\nu_{f}(p_{n}))$ is called the signature of $(\mathcal{O}_{f},\nu_{f})$. Thus the numbers appearing in the $n$-tuples are the values of the ramification function, which assign “weights” to points in $P_{f}$.
The Euler characteristic of $(\mathcal{O}_{f},\nu_{f})$ is
@@
Theorem A.[20] Let $f$ be a degree $d$ branched covering of the sphere, where $|d|>1$. Assume that there exists a completely invariant simply connected region $R$ whose boundary is locally connected. Assume moreover that it is not the case that there exists only one critical point in the boundary of $R$ that has multiplicity d1d-1 and is fixed by $f$. Then, $f$ has the rate.
We do not know if the hypothesis on local connectivity is necessary. A main ingredient in the proof of Theorem A is that $f$ extends continuously to the prime end closure of $R$. This extension of $f$ induces a map $\tilde{f}$ of the circle that turns to be a degree $d$ covering, despite of the fact that $f$ could have critical points on the boundary of the region $R$. It comes that for every positive $n$, $\tilde{f}^{n}$ has at least $|d^{n}-1|$ fixed points. As the boundary of $R$ is locally connected, to each periodic point of $\tilde{f}$ corresponds a periodic point of $f$ in the boundary of $R$. However, this correspondece is not injective, so in order to get the rate one has to understand how many different $\tilde{f}-$periodic prime ends correspond to the same point in the boundary of $R$.
@@
So, in particular, Theorem A is a topological version of the fact that complex polynomials (with connected and locally connected Julia set) have the rate. Following the polynomial analogy, the opposite situation corresponds to the case when all critical points belong to the complement of the filled Julia set. We also gave a topological version of the fact that such maps have the rate:
Theorem B.[20] Let $f$ be a degree $d$ branched covering of the sphere, where $|d|>1$. Assume that there exists a simply connected open set $U$ whose closure is disjoint from the set of critical values and such that $\overline{f^{-1}(U)}\subset U$. Then $f$ has the rate.
Going back to Theorem A, we already pointed out that local connectivity is a technical hypothesis used only to guarantee that external rays land. If one wanted to push the proof forwards, one has to consider the case where external rays land on proper subcontinua of $\partial R$ and try to fish for fixed points inside these continua. This strategy could work if for instance one could prove that the impression of periodic rays (or at least of enough of them) is contained in an injectivity region of $f$. Indeed, if that were the case, one could get fixed points of the iterates $f^{n}$ by applying some Brouwer theory locally. However, it may as well happen that the impression of every (periodic) prime end is the whole set $\partial R$, in which case this particular strategy falls appart. The question that arises then is whether or not that is possible. It is fascinating that this is not known even in the holomorphic case. Indeed, whether or not there exists a rational function with an indecomposable continuum as its Julia set is a well-known unsolved problem (see [8]). Furthermore, it was not known until recently if there existed branched coverings of the sphere of degree $d$, with $|d|>1$, having an indecomposable completely invariant continuum. In [22] such an example was given, and we will discuss it in the next section. However, this example does not have a completely invariant proper and simply-connected region. Moreover, it is not known whether or not there exists a plane branched covering of degree $d$, with $|d|>1$, having an indecomposable completely invariant continuum. In [14] the results in [8] were studied from a topological viewpoint, and were shown to be purely topological (independent of the holomorphic assumptions).
@@
As explained in the previous section, the hypothesis of local connectivity does not seem to be relevant in Theorem A and it is plausible that this hypothesis should be replaced by decomposability instead. It is then natural to try and understand the examples presenting totally invariant indecomposable continua. This turns out to be a rather intricate endeavour.
We first point out that it is well known that surface homeomorphisms admit indecomposable continua as invariant subsets. There are several methods to construct these examples, and they can even be made minimal (in particular, fixed and periodic point free). It is also well-known that the solenoid admits non-trivial connected coverings (see [7]), but the solenoid is not embeddable in the sphere. It is also known that for any integer $k$, the pseudo-circle is a $k$-fold covering space of itself (see [13] and [15]). However, the following remains unknown: does there exists a plane branched covering with a totally invariant pseudo-circle?. That is, it is not known how to extend a self-covering of the pseudo-cirle to the whole plane so that the extension is a branched covering. It is known how to extend it to a map of the annulus, which was already noted in [3]. Moreover, in [3] the rate was proven to hold for annulus maps preserving a pseudo-circle and this result was generalized to all plane separating circle-like continua.
@@
[22](“Plykin-Iglesias”). There exists a branched covering $f:S^{2}\to S^{2}$ supporting a completely invariant continuum $K$ and satisfying the following properties:
@@
Moreover -and unfortunately- this map is not Thurston equivalent to a rational map. It is a parabolic $(2,2,2,2)$ orbifold with associated torus map in the homotopy class of Anosov, which is a well-known Thurston obstruction (see Section 4 for the definitions and further references). So, even in the $C^{0}$ setting it is unknown if a map supports an indecomposable “Julia set”. More precisely, whether or not there exists a Thurston map Thurston equivalent to a rational map, having an indecomposable completely invariant continuum remains unknown. And adding to the shortcomings of our example, it is not a topological polynomial, that is, it is not a plane branched covering. Plane branched coverings were treated in Section 5, and were our main motivation to dive into dynamics of indecomposable continua. We remark that it remains unknown whether or not there exists a plane branched covering of degree $d$, with $|d|>1$, having a completely invariant indecomposable continuum.
@@
Branched coverings of the pseudo-arc were constructed in [4]. These are degree $d$ branched coverings with one critical point of multiplicity d1d-1. In [23] we provided an example of a branched covering of the Knaster continuum with a different ramification portrait, whose peculiarity is that it does not satisfy the classical Riemann-Hurwitz formula.
@@
We proved that any branched self-covering of the Knaster continuum does not verify the Riemann-Hurwitz formula, and therefore cannot be extended to a branched covering of the sphere. This result shows that studying intrinsic dynamics of coverings and branched coverings on indecomposable continua is a useful tool, even if our goal is understanding their interaction with an ambient space dynamics. It also rules out the possibility of the Knaster continuum as a totally invariant set for a branched covering of the sphere. We expect this kind of result to generalize for a wider class of indecomposable continua.
@@
This section contains a potpurri of examples that are relevant to everything discussed in this paper.

### v2

@@ 1. Introduction
(1) $\limsup\frac{1}{n}\log(\#\operatorname{Fix}(f^{n}))\geq\log(d)$
for a dynamical system $f$ of positive degree $d$ is both simple and natural. For instance, if $f:S^{1}\to S^{1}$ is continuous with degree $d$, the inequality is easily seen to be true. Indeed, the graph of a lift $\tilde{f}$$\mathbb{R}$$\to{\mathbb{R}}$ must intersect at least d1d-1 times the set $\{y=x+n,n\in{\mathbb{Z}}\}$. This gives at least $d^{n}-1$ fixed points for the map $f^{n}$. However, in contrast, one can easily construct a continuous map $f:S^{2}\to S^{2}$ with degree $d$ where inequality (1) does not hold. The map expressed in polar coordinates in ${\mathbb{R}}^{2}$ as $(r,\theta)\to(dr,d\theta)$ and extended to the sphere with $\infty\to\infty$, has degree $d$ and just two periodic points.
More interestingly, to this day there is no list of well understood obstructions for the inequality to hold for the two dimensional sphere. Moreover, in trying to understand this simple yet fundamental question of low dimensional dynamics, there are other significant open questions and problems that arise.
When an endomorphism $f$ does verify the inequality (1), we say that $f$ has the rate.
This paper provides an introductory survey of basic results obtained to date on this “rate” problem, and presents a number of related open questions and directions for further research.
To finish this introductory section I would like to add that it was Mike Shub that grew awareness around the growth rate inequality problem within the endomorphism group working in Montevideo, Uruguay, during one of his many stays in our country. We are deeply thankful for his contribution, that resulted in many years of both joy and challenge in our work.
@@ 2. $C^{1}$ scenario
We have already seen that if $f$ is a continuous map, the inequality (1) may not hold (i.e. $f$ may not have the rate). On the other hand, if $f$ is a rational map, then it has the rate (see Proposition 1 of [35]). From [36] it follows that if $f$ is $C^{1}$, then it must have infinitely many distinct periodic points, but their growth rate remains unknown. Of course this rises the question as to whether certain regularity will imply that the map has the rate. This is not the focus of the present paper. Nonetheless, we present in this section a very brief overview of the results regarding differentiable maps.
@@
The latitude preserving hypothesis means that $f$ carries each latitude into another latitude or to one of the poles. It need not be a homeomorphism from one latitude to another. Of course, it follows from Theorem 2.1 that $f$ has the rate.
@@
These sets of hypothesis (smooth and preserves some “geographical” (singular) foliation) were later considered in [13], where the authors considered what happens when one changes the “geography” and replace the assumption of smoothness by another one. More precisely, certain foliations with only one sigularity are specified and the map is required to be smooth near this “pole” in order to have the rate.
The research in [13] carried on by studying “toy models” which may help one to understand the mechanism of creating periodic points under the assumption of smoothness. See [14], where the authors confirm Shub’s conjecture for smooth maps of $S^{2}$ preserving the longitudinal foliation.
These strong hypothesis of preserving a foliation and being of class $C^{1}$ were dropped and replaced by an assumption on the homological action of the map that generalizes the former series of results:
@@
[32] Let $f:S^{2}\to S^{2}$ be a continuous map such that $\deg f=d,|d|>1$. Assume the following hypothesis hold:
@@
Let $f:S^{2}\to S^{2}$ be a degree $d$ $C^{1}$ endomorphism with $|d|>1$. If $f^{-1}(\{N,S\})\subset\{N,S\}$, then $f$ has the rate.
This result was reobtained in [1] also as a corollary of their work on the growth rate inequality for higher dimensional spheres.
A crucial remark, that is responsible for these last results and that will be exploited in the more detailed discussion to follow, is that the $C^{1}$ hypothesis is only used to guarantee that the map is a sink around its critical points. It is natural then not to impose differentiability, but rather try to understand what the real obstructions are for a sphere endomorphism to have the rate.
@@
[19] Let $f:S^{2}\to S^{2}$ be a degree $d$ map such that all its periodic orbits are isolated as invariant sets. Then, if f has no sources of degree $r$ with $|r|>1$ (this happens, in particular, if $f$ is $C^{1}$) we have that $f$ has the rate.
@@
Sphere endomorphisms are too large of a universe to handle. It is natural to focus on the simplest cases at the beginning. Other than homeomorphisms, the simplest maps one can think of, in terms of their local properties, are covering maps. A covering map $f:X\to Y$ is a continuous map such that every $y\in Y$ has an open neighborhood $V$ whose preimage is a disjoint union of open sets and each of them is mapped homeomorphically onto $V$. If, however, one is interested in dynamics, one needs $X$ to be equal to $Y$, that is, one wants a self-covering $f:X\to X$ in order to be able to iterate the map. However, the only possible covering maps $f:S^{2}\to S^{2}$ are homeomorphisms. So, if one wants to escape the invertible world, one needs to allow critical points to be present. The best generalization is the following:
@@
We will show in this section that any continuous map of the closed annulus of degree $d$ ($|d|>1$) is semi-conjugate to $z^{d}$ on $S^{1}$. Moreover, if $f:A\to A$ has degree $d$ and $K\subset A$ is compact, connected, essential (i.e. not contained in a topological disc of $A$) and invariant, then $f:|_{K}:K\to K$ is semi-conjugate to $z^{d}$ on $S^{1}$. However, this is not true for maps acting on the open annulus $A$ (as the ones arising from restricting a branched covering of the sphere with a set of two points that is totally invariant); even if they are covering maps. We refer the reader to Section 7.3 for an example of $f:A\to A$ a covering map of degree $d$ that is not semi-conjugate to $z^{d}$ acting on $S^{1}$. We denote $\tilde{A}={\mathbb{R}}\times(0,1)$ the universal cover of $A$.
@@
Another simple idea gives the rate for maps of the annulus of degree $d$, $|d|>1$: any lift $F$ will push vertical lines that are sufficiently far to the right or left strictly to the right of themselves, or strictly to the left of themselves (we are using $|d|>1$ here). So, with hypothesis guaranteeing that horizontal lines are also pushed strictly above them, or strictly below them, one can calculate the Lefschetz index of a simple closed curve formed by intersecting appropriate pairs of vertical and horizontal lines. This is formalized as follows:
@@
[22] Let $F$ be a map of the open annulus $A$ that interchanges the ends of $A$. If $|\deg(F)|>1$, then $F$ has the rate.
@@
The orbifold associated to a Thurston map $f$ is the pair $\mathcal{O}_{f}=(S^{2},\nu_{f})$. Note that $\{p\in S^{2}:\nu_{f}(p)\geq 2\}=P_{f}$ is a finite set. If we label these points $p_{1},\dots,p_{n}$ such that $2\leq\nu_{f}(p_{1})\leq\ldots\leq\nu_{f}(p_{n})$, then the $n$-tuple $(\nu_{f}(p_{1}),\ldots,\nu_{f}(p_{n}))$ is called the signature of $(S^{2},\nu_{f})$. Thus the numbers appearing in the $n$-tuples are the values of the ramification function, which assign “weights” to points in $P_{f}$.
The Euler characteristic of $(S^{2},\nu_{f})$ is
@@
Theorem 5.1.
[23] Let $f$ be a degree $d$ branched covering of the sphere, where $|d|>1$. Assume that there exists a completely invariant simply connected region $R$ whose boundary is locally connected. Assume moreover that it is not the case that there exists only one critical point in the boundary of $R$ that has multiplicity d1d-1 and is fixed by $f$. Then, $f$ has the rate.
We do not know if the hypothesis on local connectivity is necessary. A main ingredient in the proof of Theorem 5.1 is that $f$ extends continuously to the prime end closure of $R$. This extension of $f$ induces a map $\tilde{f}$ of the circle that turns to be a degree $d$ covering, despite of the fact that $f$ could have critical points on the boundary of the region $R$. It follows that for every positive $n$, $\tilde{f}^{n}$ has at least $|d^{n}-1|$ fixed points. As the boundary of $R$ is locally connected, to each periodic point of $\tilde{f}$ corresponds a periodic point of $f$ in the boundary of $R$. However, this correspondece is not injective, so in order to get the rate one has to understand how many different $\tilde{f}-$periodic prime ends correspond to the same point in the boundary of $R$.
@@
So, in particular, Theorem 5.1 is a topological version of the fact that complex polynomials (with connected and locally connected Julia set) have the rate. Following the polynomial analogy, the opposite situation corresponds to the case when all critical points belong to the complement of the filled Julia set. We also gave a topological version of the fact that such maps have the rate:
Theorem 5.2.
[23] Let $f$ be a degree $d$ branched covering of the sphere, where $|d|>1$. Assume that there exists a simply connected open set $U$ whose closure is disjoint from the set of critical values and such that $\overline{f^{-1}(U)}\subset U$. Then $f$ has the rate.
Going back to Theorem 5.1, we already pointed out that local connectivity is a technical hypothesis used only to guarantee that external rays land. If one wanted to push the proof forward, one has to consider the case where external rays land on proper subcontinua of $\partial R$ and try to fish for fixed points inside these continua. This strategy could work if for instance one could prove that the impression of periodic rays (or at least of enough of them) is contained in an injectivity region of $f$. Indeed, if that were the case, one could get fixed points of the iterates $f^{n}$ by applying some Brouwer theory locally. However, it may as well happen that the impression of every (periodic) prime end is the whole set $\partial R$, in which case this particular strategy falls appart. The question that arises then is whether or not that is possible.
It is fascinating that this is not known even in the holomorphic case. Indeed, whether or not there exists a rational function with an indecomposable continuum as its Julia set is a well-known unsolved problem (see [10]). Furthermore, it was not known until recently if there existed branched coverings of the sphere of degree $d$, with $|d|>1$, having an indecomposable completely invariant continuum. In [25] such an example was given, and we will discuss it in the next section. However, this example does not have a completely invariant proper and simply-connected region. Moreover, it is not known whether or not there exists a plane branched covering of degree $d$, with $|d|>1$, having an indecomposable completely invariant continuum.
In [16] the results in [10] were studied from a topological viewpoint, and were shown to be purely topological (independent of the holomorphic assumptions).
@@
As explained in the previous section, the hypothesis of local connectivity does not seem to be relevant in Theorem 5.1 and it is plausible that this hypothesis should be replaced by decomposability instead. It is then natural to try to understand the examples presenting totally invariant indecomposable continua. This turns out to be a rather intricate endeavour.
We first point out that it is well known that surface homeomorphisms admit indecomposable continua as invariant subsets. There are several methods to construct these examples, and they can even be made minimal; in particular, periodic point free (see, for example [17] and [3]). It is also well-known that the solenoid admits non-trivial connected coverings (see [9]), but the solenoid is not embeddable in the sphere. It is also known that for any integer $k$, the pseudo-circle is a $k$-fold covering space of itself (see [15] and [18]). However, the following remains unknown: does there exists a plane branched covering with a totally invariant pseudo-circle? That is, it is not known how to extend a self-covering of the pseudo-cirle to the whole plane so that the extension is a branched covering. It is known how to extend it to a map of the annulus, which was already noted in [5]. Moreover, in [5] the rate was proven to hold for annulus maps preserving a pseudo-circle and this result was generalized to all plane separating circle-like continua.
@@
[25] There exists a branched covering $f:S^{2}\to S^{2}$ supporting a completely invariant continuum $K$ and satisfying the following properties:
@@
Moreover -and unfortunately- this map is not Thurston equivalent to a rational map. [^3] We will not define Thurston equivalence here, but we refer the reader to Section 2.4 in [4] It is a parabolic $(2,2,2,2)$ orbifold with associated torus map in the homotopy class of Anosov, which is a well-known Thurston obstruction (see Section 4 for the definitions and further references). So, even in the $C^{0}$ setting it is unknown if a map supports an indecomposable “Julia set”. More precisely, whether or not there exists a Thurston map Thurston equivalent to a rational map, having an indecomposable completely invariant continuum remains unknown. And adding to the shortcomings of our example, it is not a topological polynomial, that is, it is not a plane branched covering. Plane branched coverings were treated in Section 5, and were our main motivation to dive into dynamics of indecomposable continua.
We remark that it remains unknown whether or not there exists a plane branched covering of degree $d$, with $|d|>1$, having a completely invariant indecomposable continuum.
@@
Branched coverings of the pseudo-arc were first constructed in [37], and constructed again and used in [6] to prove the Barge Entropy Conjecture. These are degree $d$ branched coverings with one critical point of multiplicity d1d-1. In [26] we provided an example of a branched covering of the Knaster continuum with a different ramification portrait, whose peculiarity is that it does not satisfy the classical Riemann-Hurwitz formula.
@@
We proved in [26] that any branched self-covering of the Knaster continuum does not verify the Riemann-Hurwitz formula, and therefore cannot be extended to a branched covering of the sphere. This result shows that studying intrinsic dynamics of coverings and branched coverings on indecomposable continua is a useful tool, even if our goal is understanding their interaction with an ambient space dynamics. It also rules out the possibility of the Knaster continuum as a totally invariant set for a branched covering of the sphere. We expect this kind of result to generalize for a wider class of indecomposable continua.
@@
This section contains a potpourri of examples that are relevant to everything discussed in this paper.